OpenStudy (anonymous):

Determine the number and type of solutions for –3x2 + 3x + 1= 0. 1 Rational Double 2 Irrational 2 Rational 2 Complex

OpenStudy (anonymous):

easiest just to solve and see what you get, but you can also compute \[b^2-4ac\] and do the problem that way

OpenStudy (anonymous):

\[3^2-4\times (-3)\times 1=9+12=21\] and since \(21>0\) there are two solutions irrational since 21 is not a perfect square